# TopoSPAM: Topology grounded Simulation Platform for morphogenesis and biological Active Matter

### v1

@@ Abstract
Figure 1: Modeling and simulation of biological active matter across scales. Cytoskeleton (left column): Microtubule mixtures with Kinesin-1 motor proteins exhibiting a bend instability (top panel) (2). An active polar fluid model (middle) captures the instability in the corresponding TopoSPAM Simulation (bottom panel). Cellular organoids (center column): Pancreatic organoids exhibiting spontaneous chiral rotations that are modeled with the 3D vertex model (top and middle) (3). The TopoSPAM simulation of the vertex model predicts the spontaneous chiral flows (bottom). Tissue 3D shape change (right column): Drosophila wing disc development (top, image courtesy: Natalie A. Dye) modeled via a spring-lattice model (middle). The TopoSPAM simulation of the spring-lattice model predicts the three-dimensional development of the wing disc (bottom). Notice that the models are not restricted to the scales shown here. In particular, the active polar gel model can also be used across scales.
@@ 0.1 Active polar fluids and the cytoskeleton
TopoSPAM includes a state-of-the-art method to numerically solve 2D and 3D active polar and nematic hydrodynamic problems in arbitrary geometries and with arbitrary boundary conditions (11). The hydrodynamic equations of active fluids describe the interaction between motile microscopic constituents and their surrounding fluid under energy input. These equations capture many emergent phenomena in biology, for example the dynamics of the cytoskeleton and of active tissue-scale flows (12). A mean-field description of such an out-of-equilibrium system has been derived from first principles (conservation laws, symmetries, entropy production, and the Onsager reciprocal relations) and recapitulates the resulting dynamics (13, 14, 15, 16). This model includes two continuous fields: (1) an orientation field to capture the average orientation of the microscopic constituents (e.g., cytoskeletal filaments, motor proteins) at each point in space and time; and (2) a velocity field of the emerging flows and deformations (see Methods section for the governing equations and implementation details).
@@
TopoSPAM’s active polar fluid solver has recently been used to confirm the analytically predicted 3D spontaneous flow transition of cytoskeletal suspensions (18, 11) and to recapitulate experimentally observed the out-of-plane wrinkling behavior (19) in microtubule suspensions (Fig. 3B). Here, we additionally use it to simulate an active polar fluid confined to a 3D cylindrical domain where it develops a moving spiral-shaped line defect under rotational flow (Fig. 3C). In the same geometry, a stable aster line defect induces no flow (Fig. 3D), highlighting the important role of topology in determining the morphogenetic mode. The last two cases also demonstrate the ability of TopoSPAM to solve complex active fluid problems in non-Cartesian domains.
@@ 0.2 Active vertex models and cells in tissues
Moving up the scales of biology, TopoSPAM is also capable of capturing discrete cellular behaviors inside an epithelial tissue via its 3D active vertex model framework. A potential approach to linking tissue morphogenesis with mechanistic cell biology involves “decomposing” the tissue deformation into distinct, independent cell behaviors, essentially forming its “deformation basis". Often, these cellular behaviors are intricately interconnected and carefully regulated by numerous mechano-chemical cues to govern the development of tissues and organs, ultimately shaping them into their final forms. Some common cell behaviors include individual cell elongations, T1 rearrangements, cell divisions, apoptosis, cell extrusion, and relative apical–basal surface contractions (20, 21).
One of the challenges in comprehending and establishing accurate theories of tissue morphogenesis arises from tissues’ ability to modify their internal topology, such as the neighbor network of the cells. This additional layer of complexity makes conventional theoretical frameworks, like linear elasticity and stability analysis, less applicable. To overcome this challenge, a promising approach is to work with models that effectively exploit the discrete topology of the tissue. TopoSPAM provides this through a robust implementations of 3D vertex models and spring-lattice models (22, 23).
Vertex models provide a powerful framework widely used in tissue mechanics, offering valuable insights into how individual cells’ collaborative efforts yield complex responses within a tissue (24, 25). With vertex models, we can understand how cells contribute to shaping the tissue and potentially link deformations to molecular mechanisms.
TopoSPAM also extends this to active vertex models. In contrast to many passive vertex models found in the literature (24, 22) our version also considers the active forces that the cells exert on their surroundings. In this way, TopoSPAM’s active vertex model enables us to effectively gauge whether active cellular forces or passive relaxation into energetically favorable states drives observed cell and tissue shape changes.
@@ 0.1 Polar Active Hydrodynamics
TopoSPAM includes the 2D and 3D active and passive Ericksen-Leslie equations modeling an incompressible polar fluid in the hydrodynamic limit (16, 12). This model describes the behavior of viscous active polar fluids under non-equilibrium, nonlinear hydrodynamic conditions. It relates the dynamics of two fields: (1) the macroscopic orientation field $\mathbf{p}$ with components $p_{\alpha}$ defined by the average orientation of the microscopic constituents in an infinitesimal volume; (2) the velocity field $\mathbf{v}$ with components $v_{\alpha}$ of the emerging flows and deformations. The governing Stokes equation for the force balance is
@@
The time dynamics of the fields is resolved using the adaptive Adams–Bashforth–Moulten predictor-corrector time stepping method of convergence order two. The entire active fluid simulation code is written in the OpenFPM C++ expression system (5) in only 423 lines of C++ source code that transparently parallelizes to shared- and distributed-memory computers and to Graphics Processing Units (GPUs) (11, 7).

### v2

@@ Abstract
Figure 1: Modeling and simulation of biological active matter across scales. Cytoskeleton (left column): Microtubule mixtures with Kinesin-1 motor proteins exhibiting a bend instability (top panel) (2). An active polar fluid model (middle) captures the instability in the corresponding TopoSPAM Simulation (bottom panel). Cellular organoids (center column): Pancreatic organoids exhibiting spontaneous chiral rotations that are modeled with the 3D vertex model (top and middle) (3). The TopoSPAM simulation of the vertex model predicts the spontaneous chiral flows (bottom). Tissue 3D shape change (right column): Drosophila wing disc development (top) (4) modeled via a spring-lattice model (middle). The TopoSPAM simulation of the spring-lattice model predicts the three-dimensional development of the wing disc (bottom). Notice that the models are not restricted to the scales shown here. In particular, the active polar gel model can also be used across scales.
@@ 0.1 Active polar fluids and the cytoskeleton
TopoSPAM includes a state-of-the-art method to numerically solve 2D and 3D active polar and nematic hydrodynamic problems in arbitrary geometries and with arbitrary boundary conditions (12). The hydrodynamic equations of active fluids describe the interaction between motile microscopic constituents and their surrounding fluid under energy input. These equations capture many emergent phenomena in biology, for example the dynamics of the cytoskeleton and of active tissue-scale flows (13). A mean-field description of such an out-of-equilibrium system has been derived from first principles (conservation laws, symmetries, entropy production, and the Onsager reciprocal relations) and recapitulates the resulting dynamics (14, 15, 16, 17). This model includes two continuous fields: (1) an orientation field to capture the average orientation of the microscopic constituents (e.g., cytoskeletal filaments, motor proteins) at each point in space and time; and (2) a velocity field of the emerging flows and deformations (see Methods section for the governing equations and implementation details).
@@
TopoSPAM’s active polar fluid solver has recently been used to confirm the analytically predicted 3D spontaneous flow transition of cytoskeletal suspensions (19, 12) and to recapitulate experimentally observed the out-of-plane wrinkling behavior (20) in microtubule suspensions (Fig. 3B). Here, we additionally use it to simulate an active polar fluid confined to a 3D cylindrical domain where it develops a moving spiral-shaped line defect under rotational flow (Fig. 3C). In the same geometry, a stable aster line defect induces no flow (Fig. 3D), highlighting the important role of topology in determining the morphogenetic mode. The last two cases also demonstrate the ability of TopoSPAM to solve complex active fluid problems in non-Cartesian domains.
@@ 0.2 Active vertex models and cells in tissues
Moving up the scales of biology, TopoSPAM is also capable of capturing discrete cellular behaviors inside an epithelial tissue via its 3D active vertex model framework. A potential approach to linking tissue morphogenesis with mechanistic cell biology involves “decomposing” the tissue deformation into distinct, independent cell behaviors, essentially forming its “deformation basis". Often, these cellular behaviors are intricately interconnected and carefully regulated by numerous mechano-chemical cues to govern the development of tissues and organs, ultimately shaping them into their final forms. Some common cell behaviors include individual cell elongations, T1 rearrangements, cell divisions, apoptosis, cell extrusion, and relative apical–basal surface contractions (21, 22).
One of the challenges in comprehending and establishing accurate theories of tissue morphogenesis arises from tissues’ ability to modify their internal topology, such as the neighbor network of the cells. This additional layer of complexity makes conventional theoretical frameworks, like linear elasticity and stability analysis, less applicable. To overcome this challenge, a promising approach is to work with models that effectively exploit the discrete topology of the tissue. TopoSPAM provides this through a robust implementations of 3D vertex models and spring-lattice models (23, 24).
Vertex models provide a powerful framework widely used in tissue mechanics, offering valuable insights into how individual cells’ collaborative efforts yield complex responses within a tissue (25, 26). With vertex models, we can understand how cells contribute to shaping the tissue and potentially link deformations to molecular mechanisms.
TopoSPAM also extends this to active vertex models. In contrast to many passive vertex models found in the literature (25, 23) our version also considers the active forces that the cells exert on their surroundings. In this way, TopoSPAM’s active vertex model enables us to effectively gauge whether active cellular forces or passive relaxation into energetically favorable states drives observed cell and tissue shape changes.
@@ 0.1 Polar Active Hydrodynamics
TopoSPAM includes the 2D and 3D active and passive Ericksen-Leslie equations modeling an incompressible polar fluid in the hydrodynamic limit (17, 13). This model describes the behavior of viscous active polar fluids under non-equilibrium, nonlinear hydrodynamic conditions. It relates the dynamics of two fields: (1) the macroscopic orientation field $\mathbf{p}$ with components $p_{\alpha}$ defined by the average orientation of the microscopic constituents in an infinitesimal volume; (2) the velocity field $\mathbf{v}$ with components $v_{\alpha}$ of the emerging flows and deformations. The governing Stokes equation for the force balance is
@@
The time dynamics of the fields is resolved using the adaptive Adams–Bashforth–Moulten predictor-corrector time stepping method of convergence order two. The entire active fluid simulation code is written in the OpenFPM C++ expression system (6) in only 423 lines of C++ source code that transparently parallelizes to shared- and distributed-memory computers and to Graphics Processing Units (GPUs) (12, 8).